OpenStudy (anonymous):

Which number is divisible by 3?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

36

OpenStudy (cwrw238):

there are many!!

Parth (parthkohli):

1) Add the digits. 2) If the sum of the digits is divisible by 3, then the number is too.

OpenStudy (anonymous):

any number where the sum of digits is divisible by 3

OpenStudy (anonymous):

a number is divisible by 3 if its total sum of digits in no must be divisible by 3

Parth (parthkohli):

Success in two steps!

OpenStudy (anonymous):

do you have choices?

OpenStudy (anonymous):

sum of any 3 consecutive natural number is divisible by 3

OpenStudy (anonymous):

a number is divisible by 3 if its total sum of digits in no must be divisible by 3 so 36 =3+6=9 divisible by 3 so 36 divisible by 3 also 9 divisible by 3

OpenStudy (anonymous):

hmm .. nice explanation :)

OpenStudy (anonymous):

do you have to pick out of numbers?

OpenStudy (amistre64):

all numbers are divisible by 3; the questions is, what is the result spose to be?

OpenStudy (anonymous):

when you see a number, add all the digits up and if it is divisible by 3 then the whole number is divisible by 3 for example 456 -> 4+5+6=15 15 is divisible by 3, the the whole number is Hope that helps :D